OpenStudy (anonymous):

What is P (13)? What does it mean? Explain how you could use the graph, the table, or the formula to compute P (13).

OpenStudy (anonymous):

What's the entire problem?

OpenStudy (anonymous):

Alice is paid $8 per hour in her part-time job. Let t be the amount time that she works, in hours, during the week, and let P ( t ) be her gross pay (before taxes), in dollars, for the week. Write a formula showing how her gross pay depends upon the amount of time that she works. Answer: P (t) = 8t, where t is the number of hours. What is P (13)? What does it mean? Explain how you could use the graph, the table, or the formula to compute P (13). Answer: P (13) means her gross pay before taxes. But P (13) = 104 If Alice works 9 hours and 15 minutes, what will her gross pay be? Show how you know. Express the result using function notation. If Alice works 4 hours and 50 minutes, what will her gross pay be? Show how you know. Express the result using function notation.

OpenStudy (anonymous):

P(13) is where she's worked 13 hours during her week. P(13) would be 13*8=104.

OpenStudy (anonymous):

15 min=1/4 an hour so 9.25*8 = $74 and 50 min = 5/6 an hour 4.83333333333333333333333333333333*8 = about 38 and 2/3, so $38.67

OpenStudy (anonymous):

function notation f(x) = 8x f(4.8333333(repeating) = 38.6666(repeating)

OpenStudy (anonymous):

what about the last one ???

OpenStudy (anonymous):

i just did that one

hero (hero):

P (t) = 8t P(13) = 8(13) P(9.25) = 8(9.25) P(4.83) = 8(4.83)

OpenStudy (anonymous):

yes, exactly.

hero (hero):

t = hrs